..What is the total length of 32 pieces of copper wire, if each piece is 2 3/16 inches long?

... A women worked 7 1/2 hours per day for 14 days how many hours did she work?

.... Five and one-fourth dollars are to be divided among three boys. How many dollars will each receive?
.....Three hundred fifty pounds of sugar were poured into 4 3/8 pound packages. How many packages were filled?
0 solutions
: ..What is the total length of 32 pieces of copper wire, if each piece is 2 3/16 inches long?

... A women worked 7 1/2 hours per day for 14 days how many hours did she work?

.... Five and one-fourth dollars are to be divided among three boys. How many dollars will each receive?
.....Three hundred fifty pounds of sugar were poured into 4 3/8 pound packages. How many packages were filled?
0 solutions

Answer by ankor@dixie-net.com(4535)

What is the total length of 32 pieces of copper wire, if each piece is 2 3/16 inches long?
:
32 * 23/16 = 32 * 35/16
Cancel 16 into 32:
2 * 35 = 70 inches
:
;
... A women worked 7 1/2 hours per day for 14 days how many hours did she work?
:
71/2 * 14 = 15/2
Cancel 2 into 14
7 * 15 = 105 hrs
;
:
.... Five and one-fourth dollars are to be divided among three boys. How many dollars will each receive?
:
51/4 * 1/3 = 21/4 * 1/3
Cancel 3 into 21
7/4 = 13/4
:
:
.....Three hundred fifty pounds of sugar were poured into 4 3/8 pound packages. How many packages were filled?
:
350 / 43/8 = 350 / 35/8
Invert the dividing fraction and multiply
350 * 8/35
Cancel 35 into 350
10 * 8 = 80 pkgs
..